OpenStudy (anonymous):

How do I find coordinate points that represent the x and y intercepts of the inear function y=-3x-5?

OpenStudy (amistre64):

when y= 0, what does x have to be? when x = 0 what does y have to be?

OpenStudy (anonymous):

answers could be (-5,0) and (0,-5/3), or (-5/3,0)and (0,-5), or (-5,0)and (-5/3,0), or (-5/3,0)and (-5,0)

OpenStudy (amistre64):

there is only one answer :) when y = 0 the equation looks like this: 0 = -3x -5 solve for "x" when x = 0 the equation looks like this: y = -3(0) -5 y = 0-5 y = -5 ; so when x = 0, what does "y" equal?

OpenStudy (anonymous):

1/5 ......? but why

OpenStudy (anonymous):

no, wait, that can't be right. aaarrgh.

OpenStudy (amistre64):

you write out what you are doing to solve it, and ill let you know where you are going astray..

OpenStudy (anonymous):

I don't even know where to begin That's my problem.

OpenStudy (anonymous):

anyone there?

OpenStudy (amistre64):

you begin by plugging in a value of x=0 or y=0 into the equation and solving for the other variable.

OpenStudy (amistre64):

x = 0 is usually a good be another more simplistic manner is to get the equation into the form: Ax + By = C y=-3x-5 become 3x +y = -5 3(0) + y = -5 ; y = -5 (0,-5) AND 3x + 0 = -5 3x = -5 x = -5/3...... (-5/3,0)